A closed container of cubic shape is 15 cm on a side contains four times Avogadro's number of molecules at a temperature of 20.0 degree celsius.Calculate the amount of force exerted by the gas on one of the walls of the vessel.



1
Expert's answer
2022-06-13T08:23:29-0400

Given:

"a=0.15\:\rm m"

"N=4N_A"

"T=t+273=293\:\rm K"



"F=pA=\frac{N}{N_A}\frac{RT}{V}A"

"F=\frac{N}{N_A}\frac{RT}{a^3}a^2=\frac{N}{N_A}\frac{RT}{a}"

"F=4*\frac{8.31*293}{0.15}=6.49*10^4\:\rm N"